ABC, Inc. is looking for the best option to make a key component part for its new product from three different alternatives. Alternative A would require an annual fixed cost of $300,000 and a variable cost of $40 per unit, and Alternative B would require an annual fixed cost of $200,000 and a variable cost of $20 per unit. Alternative C would require an annual fixed cost of $100,000 and a variable cost of $30 per units. To gain an insight about the capacity planning decision they make, XYZ wants to determine which alternative will work best when the customer demand varies in the future.

Determine the ranges of production volume under which each alternative becomes the best. You should include a graph to compare three alternatives and identify appropriate ranges.

Please show all your work.
Answer #1

Total cost of A = TCA = 300000 + 40*N
Total cost of B = TCB = 200000 + 20*N
Total cost of B = TCC = 100000 + 30*N

Where N = production volume

As per the graph, Alternative-A is never an optimal choice.

Alternative-C is the best when the volume is between 0 to 10,000

Alternative-B is the best when the volume is 10,000 or more.